# Compressed Momentum-based Single-Point Zero-Order Algorithm for Stochastic Distributed Nonconvex Optimization

### v1

@@ 1 Introduction
With the rapid growth of datasets and the expansion of network scales, distributed optimization has attracted significant attention due to its enhanced parallel computational capabilities and inherent robustness, where agents try to cooperatively solve a problem with information exchange only limited to immediate neighbors in the network. In many networked systems, such as power systems (mao2019finite) and wireless sensor networks (kuthadi2022optimized, lei2022distributed), each computational node is typically limited to accessing information from its immediate neighbors, driven by requirements for security and scalability.
@@
To date, a wide variety of distributed algorithms have been proposed, including distributed gradient tracking algorithms (lei2022distributed, chen2024compressed), distributed ADMM (carnevale2025admm, gholami2023admm), distributed pull-push gradient-based algorithm (nedic2025ab, dimlioglu2025communication) and so on. Additionally, a popular SGD heuristic that has proven to be crucial in many applications is the use of momentum, i.e., the use of a weighted average of past gradients instead of the current gradient (sutskever2013importance). Subsequently, xia2024momentum investigated a paradigm of momentum-based systems for nonconvex constrained optimization and distributed nonconvex optimization. li2025stochastic proposed a gradient tracking stochastic distributed optimization algorithm with adaptive momentum.
@@
However, in fields such as biochemistry, commercial applications and complex engineering systems, accessing gradients may pose significant challenges due to data privacy concerns, black-box constraints, or the high computational costs involved. To handle this problem, the gradient is approximated through random sampling and finite differences in zeroth-order (ZO) optimization. Existing ZO optimization algorithms can be divided into the following categories, namely, ZO with one-point estimator and ZO with two- and multiple-point estimator. flaxman2004online proposed the ZO algorithm with one-point estimator that queries one function value at each iteration to estimate the gradient. Nowadays, some research findings are available to integrate ZO optimization with momentum-based algorithms. For instance, huang2022accelerated studied a class of accelerated zeroth-order and first-order momentum methods for both nonconvex mini-optimization and minimax-optimization. qian2023zeroth explored a ZO proximal stochastic recursive momentum algorithm.
@@
Note that the distributed optimization algorithms require each agent to exchange information with its neighboring agents to obtain the global information. However, the communication channel often has limited bandwidth. Hence, a series of compression techniques have been developed to reduce communication overhead, including standard uniform quantizers (xu2024quantized), unbiased compressors (condat2022ef), and sparsification methods (wangni2018gradient).
@@
In this paper, motivated by the well-known momentum-based optimization algorithm and by the practical limitation that querying multiple function values is often infeasible, especially in stochastic environments, we develop a Compressed Momentum-based Single-point Zeroth-Order (CMSGD_\_1P) algorithm for stochastic distributed nonconvex optimization. The contributions are summarized as follows.
@@
  • We propose a momentum-based method for stochastic distributed nonconvex optimization. Unlike singh2021squarm, we consider a more challenging gradient-free distributed setting, where gradients are approximated via noisy single-point function evaluations. Compared with mhanna2023single, our method incorporates compressed communication, which substantially reduces the communication overhead in distributed networks.
@@
  • We establish that the proposed CMSGD_\_1P algorithm can achieve a convergence rate $\mathcal{O}(\frac{1}{\sqrt[4]{T}})$ towards the exact solution with diminishing step sizes and can achieve the sublinear convergence rate ${\mathcal{O}}(\frac{1}{{\sqrt{T}}})$ towards a neighborhood of a stationary point with fixed step sizes, where $T$ is the number of iterations.
@@
  • In contrast to mhanna2023single, which assumes that $\|x\|$ is uniformly bounded. In this paper, we analyze two related auxiliary components to avoid the need for uniform boundedness assumption on $\|x\|$.
@@ 2.3 Zeroth-order optimization
Let $F(x)=\mathbb{E}_{\xi}[f(x,\xi)]$, where $f(x,\xi)$ denotes a stochastic differentiable function. As shown in flaxman2004online, the one-point estimator can be constructed as
@@
where $x$ means the optimization variable, $\gamma>0$ denotes the parameter that sets the level of smoothing. And $u$ is a random vector that follows a symmetric distribution, $d$ means the problem dimension.
@@
where $x_{i,t}\in\mathbb{R}^{d}$ is the estimation of $x$ maintained by agent $i$ at time step $t$; $\varphi_{i,t}$ denotes the additive noise; $\gamma_{g}$ is a positive parameter called the smoothing radius; $u_{i,t}$ is a random perturbation vector generated by the agent ii.
@@
$g_{i,t}$ $=\frac{d}{\gamma_{g}}u_{i,t}\tilde{f}_{i}\big(x_{i,t}+\gamma_{g}u_{i,t},\xi_{i,t}\big)$ (6)
$=\frac{d}{\gamma_{g}}u_{i,t}(f_{i}(x_{i,t}+\gamma_{g}u_{i,t},\xi_{i,t})+\varphi_{i,t}).$
@@
The perturbation vector $u_{i,t}=(u_{i,t,1},\ldots,\\ u_{i,t,d})^{\top}\in{\mathbb{R}^{d}}$ is chosen independently by each agent $i\in\mathcal{N}$ from others and previous samples. In addition, the elements of $u_{i,t}$ are assumed i.i.d. with $\mathbb{E}(u_{i,t,d})^{2}=\sigma_{1}>0$, which implies that $\mathbb{E}[u_{i,t}u_{i,t}^{T}]=\sigma_{1}I_{d}$. Then, we further assume that there exists a constant $\sigma_{2}>0$ such that $\|u_{i,t}\|\leq\sigma_{2}.$
@@
(Proposition $3.$ of mhanna2023single) Under Assumptions 4 and 5, then, $g_{i,t}$ is a biased estimator of the agent’s gradient $\nabla{F_{i}}({x_{i,t}}),\forall i\in{\mathcal{V}}$ for every $t\geq 0$, i.e.,
@@ 2.4 Compression operators
As stated in liao2022compressed is naturally satisfied by various compressors in distributed algorithms. For example, the Randk sparsifier (beznosikov2023biased), the Topk sparsifier (zou2022downlink) and norm-sign compressor (yi2022communication).
@@ 3 Main results
In this section, we propose a Compressed Momentum-based Single-point Zeroth-Order (CMSGD_\_1P) algorithm. First, we introduce the CMSGD_\_1P algorithm and its motivation in Section 3.1. Then, we provide the supporting lemmas and establish the convergence of the CMSGD_\_1P algorithm in Section 3.2.
@@ 3.1 Algorithm Description
When the gradient information is available, in yu2019linear, the following distributed first-order momentum-based algorithm was proposed to solve (1):
$m_{i,t}$ $=\beta m_{i,{t-1}}+\nabla F_{i}(x_{i,t}),$ (10a)
@@
where $m_{i,t}$ is the local momentum of agent $i$; $\eta$ means SGD step-size; $\beta$ denotes momentum coefficient.
@@
Algorithm 1 CMSGD_1P Algorithm
1:Input: Stopping time $T$, adjacency matrix $\mathbf{W}$, and positive parameters $\gamma_{x}$, $\eta$, $\beta$, $\gamma_{g}$, $u_{i,t}$.
@@ 3.2 Supporting Lemmas
For ease of presentation, we introduce the following notations:
${{\mathbf{x}_{t}}}$ $\mathrel{\mathop{\kern 0.0pt=}\limits^{\Delta}}[{x_{1,t}^{T}},.,{x_{n,t}^{T}}]^{T}\in{\mathbb{R}^{nd}},{{\bar{\mathbf{x}}}_{t}}\mathrel{\mathop{\kern 0.0pt=}\limits^{\Delta}}[{{\bar{x}}_{t}^{T}},.,{{\bar{x}}_{t}^{T}}]^{T}\in{\mathbb{R}^{nd}},$
${{\hat{\mathbf{x}}_{t}}}$ $\mathrel{\mathop{\kern 0.0pt=}\limits^{\Delta}}[{{\hat{x}}_{1,t}^{T}},.,{{\hat{x}}_{n,t}^{T}}]^{T}\in{\mathbb{R}^{nd}},{\mathbf{m}_{t}}\mathrel{\mathop{\kern 0.0pt=}\limits^{\Delta}}[{m_{1,t}^{T}},.,{m_{n,t}^{T}}]^{T}\in{\mathbb{R}^{nd}},$
${\mathbf{g}_{t}}$ $\mathrel{\mathop{\kern 0.0pt=}\limits^{\Delta}}[{g_{1,t}^{T}},.,{g_{n,t}^{T}}]^{T}\in{\mathbb{R}^{nd}},$
$\mathcal{C}(\mathbf{x})$ $\mathrel{\mathop{\kern 0.0pt=}\limits^{\Delta}}[\mathcal{C}(x_{1})^{T},.,\mathcal{C}(x_{n})^{T}]^{T}\in{\mathbb{R}^{nd}},$
@@
To begin with, we consider the following Lyapunov candidate function.
@@
(Consensus error) Let $\mathbf{x}_{t}$} be the sequence generated by Algorithm 1, we have the following bound
@@
(Compressed error) Let $\mathbf{x}_{t}$} be the sequence generated by Algorithm 1, we have the following bound
@@
$\mathbb{E}$ $\|{\eta(\beta{\mathbf{m}_{t-1}}+{\mathbf{g}_{t-1}})}\|_{F}^{2}\leq\frac{{16{\eta^{2}}{\beta^{2}}{d^{2}}\sigma_{2}^{2}L_{{f_{1}}}^{2}}}{{(1-\beta)\gamma_{g}^{2}}}\sum\limits_{k=0}^{t-1}{{\beta^{t-1-k}}}{\chi_{k}}$ (18)
$+\frac{{16{\eta^{2}}{d^{2}}\sigma_{2}^{2}L_{{f_{1}}}^{2}}}{{\gamma_{g}^{2}}}{\chi_{t-1}}+2{\eta^{2}}{\kappa_{4}}(1+\frac{{{\beta^{2}}}}{{{{(1-\beta)}^{2}}}}),$
@@
Now, we show the convergence of CMSGD_\_1P algorithm under general nonconvex settings.
@@
Suppose Assumptions 15 hold. For the total number of iterations $T\geq\max\{{c_{1}},{c_{2}},{c_{3}},{c_{4}},{c_{5}},{c_{6}},{c_{7}}\}$, each agent runs Algorithm 1 with
${\gamma_{x}}\in(max\{{m_{1}},{m_{3}}\},min\{{m_{2}},{m_{4}}\}),$
$\gamma_{g}\geq\max\{{{\tilde{\gamma}}_{1}},{{\tilde{\gamma}}_{2}},{{\tilde{\gamma}}_{3}}\},\penalty 10000\ \eta=(1-\beta)\sqrt{\frac{n}{T}},$
where $m_{1}$$m_{4}$} and $c_{1}$$c_{7}$}, $\rho,\rho_{0},\rho_{2}$, $\rho_{3}$ and ${\tilde{\gamma}}_{1}$${\tilde{\gamma}}_{3}$} are provided in Appendix 6.3. Then,
@@
Under the same assumptions and parameters settings in Theorem 1, let ${\gamma_{g}}={T^{-\frac{1}{8}}}$ and $T\geq\max\{{c_{1}},{c_{2}},{\tilde{c}_{3}},{\tilde{c}_{4}},{\tilde{c}_{5}},{\tilde{c}_{6}},{\tilde{c}_{7}}\}$, then,
@@
where
${{\tilde{c}}_{3}}$ $={(\frac{{320n{{(1-\beta)}^{2}}{\varepsilon_{0}}{d^{2}}\sigma_{2}^{2}L_{{f_{1}}}^{2}}}{{(16+5\omega)}})^{\frac{4}{3}}},$
${\tilde{c}_{4}}$ $={(\frac{{720n{\lambda^{2}}{d^{2}}\sigma_{2}^{2}L_{{f_{1}}}^{2}{{(1-\beta)}^{2}}}}{{\omega{\delta^{2}}}})^{\frac{4}{3}}},$
${\tilde{c}_{5}}$ $={(40n{(1-\beta)^{2}}{\varepsilon_{0}}{d^{2}}\sigma_{2}^{2}L_{{f_{1}}}^{2})^{\frac{4}{3}}},$
${\tilde{c}_{6}}$ $={(\frac{{192n{{(1-\beta)}^{2}}{\omega^{3}}{d^{2}}\sigma_{2}^{2}L_{{f_{1}}}^{2}}}{{\varepsilon_{5}^{3}{\varepsilon_{6}}}})^{\frac{4}{3}}},$
${\tilde{c}_{7}}$ $={(\frac{{16n{d^{2}}\sigma_{2}^{2}L_{{f_{1}}}^{2}{{(1-\beta)}^{2}}}}{{{{(\frac{{\delta{\varepsilon_{8}}}}{{30}})}^{\frac{3}{2}}}}})^{\frac{4}{3}}}.$
@@ 4 Simulations
First, the system consists of $n=6$ agents and dimension $d=30$ that exchange information through a connected undirected communication graph. And our theoretical results via numerical experiments on a nonconvex distributed binary classification task (xie2024communication):
@@
Moreover, we adopt the unbiased 2-bit quantizer in wang2025compressed, which satisfies Definition 1. The quantization operator $\mathcal{C}(\cdot)$ takes the form
@@
We compare the Algorithm 1 against distributed stochastic gradient tracking method with a one-point gradient estimator (DSGT_\_1P) in mhanna2023single and one-point zeroth-order stochastic compressed distributed primal-dual algorithm (DZOS_\_PDA).
Fig. 1 shows that CMSGD_\_1P converges slightly slower than DSGT_\_1P but still faster than DZOS_\_PDA. Furthermore, Fig. 2 illustrates that compared to other algorithms, CMSGD_\_1P converges to the same accuracy with fewer bits. This also demonstrates the efficiency of CMSGD_\_1P.

### v2

@@ 1 Introduction
With the rapid growth of datasets and the expansion of network scales, distributed optimization has attracted significant attention due to its enhanced parallel computational capabilities and inherent robustness, where agents try to cooperatively solve a problem with information exchange only limited to immediate neighbors in the network. In many networked systems, such as power systems (Mao et al. (2019)) and wireless sensor networks (Kuthadi et al. (2022), Lei et al. (2022)), each computational node is typically limited to accessing information from its immediate neighbors, driven by requirements for security and scalability.
@@
To date, a wide variety of distributed algorithms have been proposed, including distributed gradient tracking algorithms (Lei et al. (2022), Chen et al. (2024)), distributed ADMM (Carnevale et al. (2025), Gholami et al. (2023)), distributed pull-push gradient-based algorithm (Nedić et al. (2025), Dimlioglu and Choromanska (2025)) and so on. Additionally, a popular SGD heuristic that has proven to be crucial in many applications is the use of momentum, i.e., the use of a weighted average of past gradients instead of the current gradient (Sutskever et al. (2013)). Subsequently, Xia et al. (2024) investigated a paradigm of momentum-based systems for nonconvex constrained optimization and distributed nonconvex optimization. Li et al. (2025) proposed a gradient tracking stochastic distributed optimization algorithm with adaptive momentum.
@@
However, in fields such as biochemistry, commercial applications and complex engineering systems, accessing gradients may pose significant challenges due to data privacy concerns, black-box constraints, or the high computational costs involved. To handle this problem, the gradient is approximated through random sampling and finite differences in zeroth-order (ZO) optimization. Existing ZO optimization algorithms can be divided into the following categories, namely, ZO with one-point estimator and ZO with two- and multiple-point estimator. Flaxman et al. (2004) proposed the ZO algorithm with one-point estimator that queries one function value at each iteration to estimate the gradient. Nowadays, some research findings are available to integrate ZO optimization with momentum-based algorithms. For instance, Huang et al. (2022) studied a class of accelerated zeroth-order and first-order momentum methods for both nonconvex mini-optimization and minimax-optimization. Qian and Zhao (2023) explored a ZO proximal stochastic recursive momentum algorithm.
@@
Note that the distributed optimization algorithms require each agent to exchange information with its neighboring agents to obtain the global information. However, the communication channel often has limited bandwidth. Hence, a series of compression techniques have been developed to reduce communication overhead, including standard uniform quantizers (Xu et al. (2024)), unbiased compressors (Condat et al. (2022)), and sparsification methods (Wangni et al. (2018)).
@@
In this paper, motivated by the well-known momentum-based optimization algorithm and by the practical limitation that querying multiple function values is often infeasible, especially in stochastic environments, we develop a Compressed Momentum-based Single-Point Zeroth-Order (CMSPZO) algorithm for stochastic distributed nonconvex optimization. The contributions are summarized as follows.
@@
  • We propose a momentum-based method for stochastic distributed nonconvex optimization. Unlike Singh et al. (2021), we consider a more challenging gradient-free distributed setting, where gradients are approximated via noisy single-point function evaluations. Compared with Mhanna and Assaad (2023), our method incorporates compressed communication, which substantially reduces the communication overhead in distributed networks.
@@
  • We establish that the proposed CMSPZO algorithm can achieve a convergence rate $\mathcal{O}(\frac{1}{\sqrt[4]{T}})$ towards the exact solution with diminishing step sizes and can achieve the sublinear convergence rate ${\mathcal{O}}(\frac{1}{{\sqrt{T}}})$ towards a neighborhood of a stationary point with fixed step sizes, where $T$ is the number of iterations.
@@
  • In contrast to Mhanna and Assaad (2023), which assumes that $\|x\|$ is uniformly bounded. In this paper, we analyze two related auxiliary components to avoid the need for uniform boundedness assumption on $\|x\|$.
@@ 2.3 Zeroth-order optimization
Let $F(x)=\mathbb{E}_{\xi}[f(x,\xi)]$, where $f(x,\xi)$ denotes a stochastic differentiable function. As shown in Flaxman et al. (2004), the one-point estimator can be constructed as
@@
where $x$ means the optimization variable, $\gamma>0$ denotes the parameter that sets the level of smoothing. And $u$ is a random vector that follows a symmetric distribution.
@@
where $x_{i,t}\in\mathbb{R}^{d}$ is the estimation of $x$ maintained by agent $i$ at time step $t$; $\varphi_{i,t}$ denotes the additive noise at time step $t$; $\gamma_{g}$ is a positive parameter called the smoothing radius; $u_{i,t}$ is a random perturbation vector generated by the agent $i$ at time step $t$.
@@
$g_{i,t}$ $=\frac{d}{\gamma_{g}}u_{i,t}\tilde{f}_{i}\big(x_{i,t}+\gamma_{g}u_{i,t},\xi_{i,t}\big)$
$=\frac{d}{\gamma_{g}}u_{i,t}(f_{i}(x_{i,t}+\gamma_{g}u_{i,t},\xi_{i,t})+\varphi_{i,t}).$ (6)
@@
The perturbation vector $u_{i,t}=(u_{i,t,1},\ldots,\\ u_{i,t,d})^{\top}\in{\mathbb{R}^{d}}$ is chosen independently by each agent $i\in\mathcal{N}$ from others and previous samples. In addition, the elements of $u_{i,t}$ are assumed i.i.d. with $\mathbb{E}(u_{i,t,d})^{2}=\sigma_{1}>0$, which implies that $\mathbb{E}[u_{i,t}u_{i,t}^{T}]=\sigma_{1}I_{d}$. Then, we assume that there exists a constant $\sigma_{2}>0$ such that $\|u_{i,t}\|\leq\sigma_{2}.$
@@
(Proposition $3.$ of Mhanna and Assaad (2023)) Under Assumptions 4 and 5, then, $g_{i,t}$ is a biased estimator of the agent’s gradient $\nabla{F_{i}}({x_{i,t}}),\forall i\in{\mathcal{V}}$ for every $t\geq 0$, i.e.,
@@ 2.4 Compression operators
As stated in Liao et al. (2022) is naturally satisfied by various compressors in distributed algorithms. For example, the Randk sparsifier (Beznosikov et al. (2023)), the Topk sparsifier (Zou et al. (2022)) and norm-sign compressor (Yi et al. (2022)).
@@ 3 Main results
In this section, we propose a Compressed Momentum-based Single-Point Zeroth-Order (CMSPZO) algorithm. First, we introduce the CMSPZO algorithm and its motivation in Section 3.1. Then, we provide the supporting lemmas and establish the convergence of the CMSPZO algorithm in Section 3.2.
@@ 3.1 Algorithm Description
When the gradient information is available, in Yu et al. (2019), the following distributed first-order momentum-based algorithm was proposed to solve (1):
$m_{i,t}$ $=\beta m_{i,{t-1}}+\nabla F_{i}(x_{i,t-1},\xi_{i,t-1}),$ (10a)
@@
where $m_{i,t}$ is the local momentum of agent $i$ at time step $t$; $\eta$ means stochastic gradient descent (SGD) step-size; $\beta\in[0,1)$ denotes momentum coefficient.
@@
Algorithm 1 CMSPZO Algorithm
1:Input: Stopping time $T$, adjacency matrix $\mathbf{W}$, and positive parameters $\gamma_{x}$, $\eta$, $\beta$, $\gamma_{g}$.
@@ 3.2 Supporting Lemmas
For the sake of analysis, we introduce the following notations:
${{\mathbf{x}_{t}}}$ $\mathrel{\mathop{\kern 0.0pt=}\limits^{\Delta}}[{x_{1,t}^{T}},.,{x_{n,t}^{T}}]^{T},\penalty 10000\ \penalty 10000\ {{\bar{\mathbf{x}}}_{t}}\mathrel{\mathop{\kern 0.0pt=}\limits^{\Delta}}[{{\bar{x}}_{t}^{T}},.,{{\bar{x}}_{t}^{T}}]^{T},$
${{\hat{\mathbf{x}}_{t}}}$ $\mathrel{\mathop{\kern 0.0pt=}\limits^{\Delta}}[{{\hat{x}}_{1,t}^{T}},.,{{\hat{x}}_{n,t}^{T}}]^{T},\penalty 10000\ \penalty 10000\ {\mathbf{m}_{t}}\mathrel{\mathop{\kern 0.0pt=}\limits^{\Delta}}[{m_{1,t}^{T}},.,{m_{n,t}^{T}}]^{T},$
${\mathbf{g}_{t}}$ $\mathrel{\mathop{\kern 0.0pt=}\limits^{\Delta}}[{g_{1,t}^{T}},.,{g_{n,t}^{T}}]^{T},\penalty 10000\ \penalty 10000\ \mathcal{C}(\mathbf{x})\mathrel{\mathop{\kern 0.0pt=}\limits^{\Delta}}[\mathcal{C}(x_{1})^{T},.,\mathcal{C}(x_{n})^{T}]^{T},$
@@
To begin with, we consider the following Lyapunov candidate function
@@
(Consensus error) Let $\{\mathbf{x}_{t}\}$ be the sequence generated by Algorithm 1, we have the following bound
@@
(Compressed error) Let $\{\mathbf{x}_{t}\}$ be the sequence generated by Algorithm 1, we have the following bound
@@
$\mathbb{E}$ $\|{\eta(\beta{\mathbf{m}_{t-1}}+{\mathbf{g}_{t-1}})}\|_{F}^{2}\leq\frac{{16{\eta^{2}}{\beta^{2}}{d^{2}}\sigma_{2}^{2}L_{{f_{1}}}^{2}}}{{(1-\beta)\gamma_{g}^{2}}}\sum\limits_{k=0}^{t-1}{{\beta^{t-1-k}}}{\chi_{k}}$
$+\frac{{16{\eta^{2}}{d^{2}}\sigma_{2}^{2}L_{{f_{1}}}^{2}}}{{\gamma_{g}^{2}}}{\chi_{t-1}}+2{\eta^{2}}{\kappa_{4}}(1+\frac{{{\beta^{2}}}}{{{{(1-\beta)}^{2}}}}),$ (18)
@@
Now, we show the convergence of CMSPZO algorithm under general nonconvex settings.
@@
Suppose Assumptions 15 hold. For the total number of iterations $T>{\tilde{\varepsilon}_{8}}$, each agent runs Algorithm 1 with
${m_{2}}<{\gamma_{x}}<{m_{1}},\penalty 10000\ \eta=(1-\beta)\sqrt{\frac{n}{T}},$
where $m_{1}$}, $m_{2}$} and $\tilde{\varepsilon}_{8}$} are provided in Appendix 6.4. Then,
@@
Under the same assumptions and parameters settings in Theorem 1, let ${\gamma_{g}}={T^{-\frac{1}{8}}}$ and $T>{\tilde{\varepsilon}_{9}}$, then,
@@
where $\tilde{\varepsilon}_{9}$} can be found in Appendix 6.5.
Proof: The detailed proof is given in Appendix 6.5.
@@ 4 Simulations
First, the system consists of $n=6$ agents and dimension $d=30$ that exchange information through a connected undirected communication graph. And our theoretical results via numerical experiments on a nonconvex distributed binary classification task (Xie et al. (2024)):
@@
Moreover, we adopt the unbiased 2-bit quantizer in Wang et al. (2025), which satisfies Definition 1. The quantization operator $\mathcal{C}(\cdot)$ takes the form
@@
We compare the Algorithm 1 against distributed stochastic gradient tracking method with a one-point gradient estimator (DSGT_\_1P) in Mhanna and Assaad (2023) and one-point zeroth-order stochastic compressed distributed primal-dual algorithm (DZOS_\_PDA).
Fig. 1 shows that CMSPZO converges slightly slower than DSGT_\_1P but still faster than DZOS_\_PDA. Furthermore, Fig. 2 illustrates that compared to other algorithms, CMSPZO converges to the same accuracy with fewer bits. This also demonstrates the efficiency of CMSPZO.